# BugCraft: End-to-End Crash Bug Reproduction Using LLM Agents in Minecraft

### v1

@@ 3.2.1. Preprocessing
3.2.2. Step Synthesizer
Figure 4. Step Synthesizer Component
The Step Synthesizer component seen in Figure 4 is responsible for converting the unstructured text of bug reports into high-quality S2R steps, augmented with relevant external information. For the entire step synthesizing process, we use the same MLLM (Multimodal Large Language Model), which we call the Step LLM. Further details regarding the prompts utilized in the Step Synthesizer are provided in the reproduction package. This process occurs in the following stages:
@@ 3.2.2. Step Synthesizer
Knowledge Augmentation: Prior to generating the S2R steps, we aim to identify relevant wiki pages. This process begins by prompting the Step LLM with the bug report to perform named entity recognition, resulting in a list of entities. Fuzzy matching is then employed to compare this list against wiki page titles, yielding a set of potentially relevant wiki page titles. Then, the Step LLM selects the most relevant titles from the set. In addition to these specifically identified pages, we incorporate generally relevant pages, such as those pertaining to the bug report’s version, general game mechanics, and commands. Once all titles are selected, the corresponding pages are fetched, and individual analysis trajectories are generated for each page. Inspired by the self-reasoning paper (xia2024improvingretrievalaugmentedlanguage, ), which generates a reasoning trajectory for each retrieved chunk, we adopt a similar approach to analyze each page. An analysis trajectory is a step-by-step analysis generated by the Step LLM that examines the relevance of a specific external information source (e.g., a wiki page) to the bug report, highlighting key details and their potential implications for bug reproduction. This is crucial for handling the potential length of the pages and ensuring that only relevant information is cited. The LLM is also instructed to disregard any page deemed irrelevant. The resulting analysis trajectories, representing the relevant information extracted from the wiki pages, are appended to the end of every LLM call unless otherwise specified.
Initial S2R Generation: Leveraging the bug report text, comments, and curated web content, we generate an initial S2R prompting the Step LLM. This S2R is crafted to address the most critical aspects of bug reproduction while simultaneously synthesizing the unstructured text from the bug report and comments into a single, structured list. Furthermore, the use of Step LLM during this generation process enhances the S2R’s effectiveness by enabling the creation of more readily executable steps by eliminating ambiguity and promoting the use of easier and more reliable alternate steps compared to the original bug report.
@@
S2R Enhancement: Following the initial one-shot generation of the S2R, we subject it to a two-stage critique process, each stage utilizing a separate Step LLM call. This critique specifically targets inconsistencies within the plan and the predicted behaviors of in-game entities (mobs) [^10] https://minecraft.wiki/w/Mob. The initial S2R may contain errors regarding common mob behaviors, such as targeting, movement, and other related factors. To address this, we employ few-shot (tombrownfewshot, ) examples to guide the Step LLM in generating suggestions for improvement. Subsequently, we utilize two additional Step LLM calls to rewrite the S2R, placing emphasis on incorporating the generated suggestions. This iterative refinement process (madaan2023selfrefineiterativerefinementselffeedback, ) ensures the final S2R is comprehensive, internally consistent, and accurately reflects the expected behaviors of other agents in the game, such as mobs.
@@
Step Clustering: The finalized S2R is then used as input for a Step LLM prompt that instructs it to cluster the steps. This process empirically yields 2-4 clusters, each with a descriptive title and a list of its constituent steps. This clustering improves the readability and digestibility of the S2R. Following cluster generation, a critique is performed to identify any misplaced steps. Based on this critique, steps are reordered as needed. The outcome of this process is a refined and finalized S2R, optimized for clarity and downstream tasks. For instance, in MC-276621, our system effectively integrated the initial bug report and subsequent comment, enhancing the steps to reproduce with relevant commands for easier replication (Figure 5).
Figure 5. Initial MC-276621 Report and Our Output
@@
3.2.3. Action Model
Figure 6. Action Model Component
The Action Model component, illustrated in Figure 6, is responsible for translating the refined S2R into executable scripts. These scripts interact with the game interface, encompassing both the main menu and in-game elements. To achieve this, we developed a custom macro API, as existing solutions lacked a unified approach for handling both GUI and in-game interactions. Furthermore, the development of a custom macro API enabled the inclusion of game-specific methods tailored for frequently used interactions. To enhance the LLM’s ability to identify and interact with specific GUI elements, we integrated OmniParser (lu2024omniparserpurevisionbased, ). OmniParser is, as of November 2024, a state-of-the-art model framework for parsing UI screenshots into structured elements. It excels at reliably identifying interactable icons and texts within the user interface and understanding the semantics of various elements in a screenshot while providing annotations.
3.2.3.1 Base Macro API
@@
command(instruction: String): Executes Minecraft commands. Commands are restricted to non-menu contexts. This is the shortcut for the LLM to open the chat, write the command, and then send the command message.
@@
3.2.3.2 Action Model Framework
@@
For the Action Model loop, we implement a logic similar to ReAct (yao2023react, ). Further details regarding the prompts utilized in the Action Model are provided in the reproduction package. For the entire action-taking process, we use the same MLLM, which we call the Action LLM. To execute each S2R, the Action Model follows this loop:
@@
Thought & Action Generation: Prior to executing an action, the Action LLM is prompted to generate both a ”thought,” outlining its intended course of action, and the corresponding action itself in a single step. This prompt is supplied with the current trajectory (comprising previous thoughts, actions, and feedback), the annotated table generated by OmniParser, the corresponding game screenshot, and relevant examples. The combination of visual input from the game image and the structured information from the annotation table facilitates the Action LLM’s ability to identify and interact with specific game elements.
Action Verification: Following the generation of the thought and action, the Action LLM is prompted to verify whether the proposed thought and action are correct (madaan2023selfrefineiterativerefinementselffeedback, ) and conducive to progressing the game state. This verification step acts as a crucial safeguard against errors and infinite loops, where the agent might repeatedly execute the same ineffective action. If the verification fails, the generated feedback, indicating the reason for failure, is used to prompt the generation of a new thought and action pair. This revised pair is then passed to the Macro API for execution.
@@
3.2.4. S2R & Action Model Implementation Details
We implement the framework using Python. We use LangChain [^11] https://www.langchain.com/, to implement the LLM calling logic. For all LLM calls, we keep the temperature at 0 to maintain maximum determinism. For the Action Model, we only keep the last 25 iterations in Action LLM’s context to prevent context length overflow and set the iteration limit as 30. Moreover, while formatting the trajectory iterations, instead of putting the commands, we put verbal representations of what that command means. We do this to increase the interpretability of the generated actions for the LLM.
3.3. Evaluation Setup
To evaluate BugCraft, we pose two research questions about its components, considering real-world usage:
@@ 3.3. Evaluation Setup
To address RQ1 and RQ2, we conducted experiments using a curated subset of our crash bug dataset. This subset comprises 100 bug reports, carefully selected to represent a variety of scenarios that require interaction with both GUI and in-game world elements, as mentioned in section 3.1.1. The selection of this refined dataset was driven by the resource-intensive nature of running full evaluations, making it impractical to test all selected bug reports within the scope of this study.
3.3.1. Evaluation of the Step Synthesizer
To evaluate the accuracy of the step clusters generated by the Step Synthesizer, we manually reviewed each S2R. This involved attempting to reproduce the bug in Minecraft by strictly following the generated steps. If ambiguities or obstacles were encountered, we first consulted the original bug report for clarification. If the original report was insufficient, we referred to the final version of the bug report on Mojira and other relevant online resources (e.g., Minecraft Wiki) to understand the intended reproduction process.
Each step cluster was then labeled as either ”True”, ”Faulty”, or ”Irreproducible”.
@@ 3.3.1. Evaluation of the Step Synthesizer
  • Irreproducible: The bug cannot be reproduced due to reasons unrelated to the Step Synthesizer’s output. Examples include dependence on outdated versions, specific hardware, or fundamental impossibility.
To ensure the reliability of our labeling, particularly the distinction between ”Faulty” and ”Irreproducible,” we conducted an inter-rater agreement analysis. Prior to the main labeling task, we jointly inspected the results of a test run to establish the labeling categories and refine our understanding of the criteria. Subsequently, we independently labeled the 100 bug reports in our run for the step synthesizer. Inter-rater agreement was measured using both percentage agreement and Cohen’s Kappa (McHugh_2012cohen, ). The results of this analysis, including a confusion matrix, are presented 4.2.1.
@@
  1. (1)
    Wrong Command: An S2R is categorized here if it includes commands that are not recognized or parseable by Minecraft. This often occurs during the S2R enhancement phase, where the LLM may rely on its pre-trained data, which might not encompass newer command syntax, resulting in the substitution of older, invalid command formats.
  2. (2)
    Missing Step: An S2R falls under this category when it lacks a critical step, rendering the subsequent steps implausible. This issue was more prevalent with bug reports from newer Minecraft versions. This could be attributed to the increasing complexity of the game with each update and the training data cutoff date of GPT-4o (October 2023)(openai2024gpt4ocard, ), potentially leading to an incomplete understanding of the latest game mechanics.
  3. (3)
    Logic Error: This category includes steps that are impossible, significantly hinder reproduction, or are internally contradictory within the plan itself For example, simultaneously setting the time to day and spawning hostile mobs (which would burn in daylight), or attempting actions restricted by an open inventory screen.
To ensure the reliability of our labeling, we conducted an inter-rater agreement analysis. Prior to the main labeling task, we jointly inspected the results of a test run to establish the labeling categories and refine our understanding of the criteria. Subsequently, we independently labeled the 100 bug reports in our main dataset. Inter-rater agreement was measured using both percentage agreement and Cohen’s Kappa. The results of this analysis are presented in  4.2.
3.3.2. Evaluation of the Action Model
To evaluate the Action Model’s ability to reproduce bugs, we conducted a comprehensive analysis of its performance on the subset of 86 bug reports that were deemed reproducible during the Step Synthesizer evaluation (Section 3.3.1). Each reproducible report was processed end-to-end, and the Action Model’s success or failure in reproducing the bug was recorded.
3.3.3. Determining Action Model Success or Failure
@@
3.3.4. Failure Analysis and Inter-Rater Agreement
To understand the reasons behind failures, we performed a detailed examination of each unsuccessful attempt. This involved identifying recurring patterns and common failure modes, which informed the development of a structured categorization rubric. Just like we did with the Step Synthesizer rubric, we collaboratively analyzed the results of the same test run to define the failure categories, refine our understanding of the criteria, and establish a shared understanding of potential edge cases in applying the rubric. We also used our understanding of common agent mistakes to enhance the rubric. The results of this analysis, are presented in Section 4.3.
3.3.4.1 Action Model Grading Rubric
@@ A. On Step Synthesizer Success
      1. (i)
        Stuck in Loop: The agent enters a repetitive loop, hindering progress.
        1. (A)
          Menu Loop: The agent is trapped within the game’s menu system, particularly the ”Create World” screen in newer versions.
        2. (B)
          Command Loop: The agent repeatedly attempts to execute the same incorrect or unparseable command.
        3. (C)
          Death Loop: The agent continuously dies due to in-game factors, such as hostile mobs, preventing it from acting.
      2. (ii)
@@ B. On Step Synthesizer Failure
4. Results
This section presents the results of our evaluation of BugCraft, focusing on the two main components: the Step Synthesizer and the Action Model. We first analyze the Step Synthesizer’s ability to generate accurate bug reproduction plans, followed by an examination of the Action Model’s performance in executing these plans within the complete end-to-end pipeline.
4.1. Experimental Setup
Our experiments were conducted on a high-performance machine running Windows 11, equipped with 32GB of RAM, an Intel i9-14900HX CPU, and an NVIDIA RTX 4090 Mobile GPU.
We used OpenAI’s GPT-4o (specifically, the gpt-4o-2024-11-20 version) as the underlying language model for both the Step Synthesizer and Action Model components. We chose this model due to its state-of-the-art performance and widespread adoption (openai2024gpt4ocard, ). The temperature parameter for all LLM calls was set to 0 to ensure deterministic and consistent outputs(renze2024effectsamplingtemperatureproblem, ).
The complete testing process for the 100 bug reports took approximately 36 hours. This duration is attributed to the inherently sequential nature of testing each bug report, as parallelization was not feasible. Individual runs could take up to 30 minutes, depending on the complexity of the bug and the actions required for reproduction. [^12] https://minecraft.wiki/w/Java_Edition_19w13b
4.2. Step Synthesizer Analysis
Table 1. Step Synthesizer Planning Correction Analysis
@@ 4.2. Step Synthesizer Analysis
Table 1 summarizes the results of our Step Synthesizer evaluation, conducted according to the methodology outlined in Section 3.3.1. As detailed in our evaluation setup, we manually tested the generated step clusters by attempting to reproduce the bugs in the corresponding Minecraft game versions. The step clusters were categorized as ”True” if they could successfully lead to bug reproduction, ”Faulty” if they contained errors preventing reproduction, or ”Irreproducible” if the bug itself could not be reproduced even by using the initial BR, which is not a fault of the step synthesizer.
The key findings, as shown in the table, are that 66.28% of the generated step clusters were deemed ”True”, indicating a reasonable level of accuracy in the Step Synthesizer’s plan generation. However, 33.72% of the plans were categorized as ”Faulty.” Further analysis of these faulty plans revealed that the most common error was ”Wrong Command” (48.28%), followed by ”Missing Step” (34.49%) and ”Logic Error” (31.03%). Moreover, we classified 14 bug reports as ”Irreproducible,” due to factors such as dependence on specific hardware configurations or errors in the original bug report, underscoring the complexities of bug reproduction.
4.2.1. Inter-Rater Agreement for Step Synthesizer

### v2

@@ III-B1 Preprocessing
III-B2 Step Synthesizer
Figure 4: Step Synthesizer Component
As shown in Figure 4, the Step Synthesizer component is specifically designed to handle the ambiguity and incompleteness inherent in many bug reports. Its core function is to transform the unstructured, and potentially unclear, textual descriptions into high-quality S2R steps. This process includes augmenting the steps with relevant external information to fill potential gaps. We employ a single VLM (Vision Large Language Model), referred to as the Step LLM, for the entire synthesizing process. The prompts used to guide the Step LLM are detailed in the reproduction package. This conversion happens through the following stages:
@@ III-B2 Step Synthesizer
Knowledge Augmentation: Prior to generating the S2R steps, we aim to identify relevant wiki pages to employ retrieval augmented generation (RAG) [24]. This process begins by prompting the Step LLM with the bug report to perform named entity recognition, resulting in a list of entities. Fuzzy matching is then employed to compare this list against wiki page titles, yielding a set of relevant wiki page titles. Then, the Step LLM selects the most relevant titles from the set. In addition to these specifically identified pages, we incorporate generally relevant pages, such as those pertaining to the bug report’s version, general game mechanics, and commands. Once all titles are selected, the corresponding pages are fetched, and individual analysis trajectories are generated for each page. Inspired by the self-reasoning paper [25], which generates a reasoning trajectory for each retrieved chunk, we adopt a similar approach to analyze each page. A reasoning trajectory is a step-by-step analysis generated by the Step LLM that examines the relevance of a specific external information source (e.g., a wiki page) to the bug report, highlighting key details and their potential implications for bug reproduction. This is crucial for handling the potential length of the pages and ensuring that only relevant information is cited. The LLM is also instructed to disregard any page deemed irrelevant. The resulting analysis trajectories, representing the relevant information extracted from the wiki pages, are appended to the end of every LLM call after S2R generation.
Initial S2R Generation: Leveraging the bug report, comments, and web content, the Step LLM acts as a creative problem-solver to generate the initial S2R. It deconstructs the reporter’s description to find the most reliable path to reproduction. A core part of this process is generating creative solutions that leverage Minecraft’s command system to bypass difficult physical tasks. For example, instead of a step requiring the agent to physically navigate to a specific location and aim correctly to place a block, the synthesizer will generate a /setblock command. This strategic substitution is designed to minimize reliance on challenging spatial reasoning, thereby creating a clearer and more reproducible set of instructions that dramatically enhances the effectiveness of the Action Model.
@@
S2R Refinement: Following the initial one-shot generation of the S2R, we subject it to a two-stage critique process, each stage utilizing a separate Step LLM call. This critique specifically targets inconsistencies within the plan and the predicted behaviors of in-game entities (mobs) [^9] https://minecraft.wiki/w/Mob. The initial S2R may contain errors regarding common mob behaviors, such as targeting, movement, and other related factors. To address this, we employ few-shot [22] examples to guide the Step LLM in generating suggestions for improvement. Subsequently, we utilize two additional Step LLM calls to rewrite the S2R, placing emphasis on incorporating the generated suggestions. This iterative refinement process [26] ensures the final S2R is comprehensive, internally consistent, and accurately reflects the expected behaviors of other agents in the game, such as mobs.
@@
Step Clustering: The finalized S2R is then used as input for a Step LLM prompt that instructs it to cluster the steps. This process empirically yields 2-4 clusters, each with a descriptive title and a list of its constituent steps. This clustering improves the readability and digestibility of the S2R. Following cluster generation, a critique is performed to identify any misplaced steps. Based on this critique, steps are reordered as needed. The outcome of this process is a refined and finalized S2R, optimized for clarity and downstream tasks. For instance, in MC-276621, our system effectively integrated the initial bug report and subsequent comment, enhancing the S2R with relevant commands for easier replication (Figure 5).
@@
Figure 5: Initial MC-276621 Report and Our Output
III-B3 Action Model
Figure 6: Action Model Component
The Action Model component, illustrated in Figure 6, is responsible for translating the refined S2R into executable scripts. These scripts interact with the game interface, encompassing both the main menu and in-game elements. To achieve this, we developed a custom macro API, as existing solutions lacked a unified approach for handling both GUI and in-game interactions. Furthermore, the development of a custom macro API enabled the inclusion of game-specific methods tailored for frequently used interactions. To enhance the LLM’s ability to identify and interact with specific GUI elements, we integrated OmniParser [27]. OmniParser is a state-of-the-art open-weights model framework for parsing UI screenshots into structured elements. It excels at reliably identifying interactable icons and texts within the user interface and understanding the semantics of various elements in a screenshot while providing annotations.
III-B3a Base Macro API
@@ III-B3 Action Model
  • command(instruction: String): Executes Minecraft commands. Commands are restricted to non-menu contexts by the game engine. This is the shortcut for the LLM to open the chat, write the command, and then send the command message.
@@
III-B3b Action Model Framework
@@
For the Action Model loop, we implement a logic similar to ReAct [21]. Further details regarding the prompts utilized in the Action Model are provided in the reproduction package. For the entire action-taking process, we use the same VLM, which we call the Action LLM. To execute each S2R, the Action Model loop is as follows:
@@
Thought & Action Generation: Prior to executing an action, the Action LLM is prompted to generate both a ”thought,” outlining its intended course of action, and the corresponding action itself in a single step. This prompt is supplied with the current trajectory (comprising previous thoughts, actions, and feedback), the annotated table generated by OmniParser, the corresponding game screenshot, and few shot examples to improve performance [22]. The combination of visual input from the game image and the structured information from the annotation table facilitates the Action LLM’s ability to identify and interact with specific game elements.
Action Verification: Following the generation of the thought and action, the Action LLM is prompted to verify whether the proposed thought and action are correct [26] and conducive to progressing the game state. This verification step acts as a crucial safeguard against errors and infinite loops, where the agent might repeatedly execute the same ineffective action. If the verification fails, the generated feedback, indicating the reason for failure, is used to prompt the generation of a new thought and action pair. This revised pair is then passed to the Macro API for execution.
@@
III-B4 Implementation Details
We implement the framework using Python. We use LangChain [^10] https://www.langchain.com/, to implement the LLM calling logic. For the Action Model, we only keep the last 25 iterations in Action LLM’s context to prevent context length overflow and set the iteration limit as 30 for our termination case. Moreover, while formatting the trajectory iterations, instead of putting the commands, we put verbal representations of what that command means. We do this to increase the interpretability of the generated actions for the LLM.
For UI-Tars-1.5-7B, we adapt the implementation from OSWorld benchmark[28] as it was done by the model’s authors. We host the model on RunPod111111runpod.io, using VLLM[29]. For OpenAI-CUA, we again adapt OSWorld’s implementation while following the official example on how to use OpenAI-CUA given by OpenAI [^12] https://github.com/openai/openai-cua-sample-app for guidance. We run both models using the inital BRs with 30 iterations.
III-C Evaluation Setup
To evaluate BugCraft, we pose three research questions about its components, considering real-world usage and comparative performance:
@@ III-C Evaluation Setup
RQ3: How does BugCraft compare to existing general GUI agents and specialized computer-use models?
To address RQ1 and RQ2, we conducted experiments using a curated subset of our crash bug dataset, which we call the BugCraft-Bench. For RQ3, we evaluated BugCraft against UI-TARS-1.5-7B and OpenAI’s Computer Use Agent on the same dataset.
III-C1 Evaluation of the Step Synthesizer
To evaluate the accuracy of the step clusters generated by the Step Synthesizer, we manually reviewed each S2R. This involved attempting to manually reproduce the bug in Minecraft by strictly following the generated steps. If ambiguities or obstacles were encountered, we first consulted the original bug report for clarification. If the original report was insufficient, we referred to the final version of the bug report on Mojira and other relevant online resources (e.g., Minecraft Wiki) to understand the intended reproduction process. Each step cluster was then labeled as either ”True”, ”Faulty”, or ”Irreproducible”.
@@ III-C1 Evaluation of the Step Synthesizer
  • Irreproducible: The bug cannot be reproduced due to reasons unrelated to the Step Synthesizer’s output. Examples include dependence on outdated versions, specific hardware, fundamental impossibility or the bug not causing a crash when reproduced.
To ensure the reliability of our labeling, particularly the distinction between ”Faulty” and ”Irreproducible,” we conducted an inter-rater agreement analysis. Prior to the main labeling task, we jointly inspected the results of a test run to establish the labeling categories and refine our understanding of the criteria. Subsequently, we independently labeled the 100 bug reports in our run for the step synthesizer. Inter-rater agreement was measured using both percentage agreement and Cohen’s Kappa [30]. The results of this analysis, including a confusion matrix, are presented in section IV-B1.
@@
  1. 1.
    Wrong Command: An S2R is categorized here if it includes commands that are not recognized or parseable by Minecraft.
  2. 2.
    Missing Step: An S2R falls under this category when it lacks a critical step, rendering the subsequent steps implausible.
  3. 3.
    Logic Error: This category includes steps that are impossible, significantly hinder reproduction, or are internally contradictory within the plan itself. For example, simultaneously setting the time to day and spawning hostile mobs (which would burn in daylight), or attempting actions restricted by an open inventory screen.
To ensure the reliability of our labeling, we conducted an inter-rater agreement analysis. Prior to the main labeling task, we jointly inspected the results of a test run to establish the labeling categories and refine our understanding of the criteria. Subsequently, we independently labeled the 100 bug reports in our main dataset. Inter-rater agreement was measured using both percentage agreement and Cohen’s Kappa. The results of this analysis are presented in section  IV-B.
III-C2 Evaluation of the Action Model
To evaluate the Action Model’s ability to reproduce bugs, we conducted a comprehensive analysis of its performance on the subset of 86 bug reports that were deemed reproducible during the Step Synthesizer evaluation (Section III-C1). Each reproducible report was processed end-to-end, and the Action Model’s success or failure in reproducing the bug was recorded.
III-C3 Determining Action Model Success or Failure
@@
III-C4 Failure Analysis and Inter-Rater Agreement
To understand the reasons behind failures, we performed a detailed examination of each unsuccessful attempt. This involved identifying recurring patterns and common failure modes, which led to the development of a structured categorization rubric. Just like we did with the Step Synthesizer rubric, we collaboratively analyzed the results of the same test run to define the failure categories, refine our understanding of the criteria, and establish a shared understanding of potential edge cases in applying the rubric. We also used our understanding of common agent mistakes to enhance the rubric. The results of this analysis, are presented in section IV-C.
@@ A. On Step Synthesizer Success
      1. i.
        Stuck in Loop: The agent enters a repetitive loop, hindering progress. (e.g., Menu Loop, Command Loop, Death Loop).
      2. ii.
@@ B. On Step Synthesizer Failure
IV Results
This section evaluates BugCraft’s Step Synthesizer and Action Model end-to-end, then benchmarks baseline models and an additional model on our framework.
IV-A Experimental Setup
Our experiments were conducted on a high-performance machine running Windows 11, equipped with 32GB of RAM, an Intel i9-14900HX CPU, and an NVIDIA RTX 4090 Mobile GPU. We used OpenAI’s GPT-4o (specifically, the gpt-4o-2024-11-20 snapshot) as the underlying language model for both the Step Synthesizer and Action Model components; we chose this model for its state-of-the-art performance and widespread adoption [31]. To demonstrate that our framework is model-agnostic we also evaluated the identical framework with GPT-4.1 (gpt-4.1-2025-04-14), keeping every other hyper-parameter unchanged. All LLM calls in our framework were executed with temperature = 0 to ensure deterministic outputs [32]. OmniParser served as our annotation backbone owing to its superior accuracy over comparable tools, and baselines (OpenAI CUA and UI-TARS-1.5-7B) are detailed in Section IV-D.
IV-B Step Synthesizer Analysis
TABLE I: Step Synthesizer Planning Correction Analysis
@@ IV-B Step Synthesizer Analysis
Table I summarizes the Step Synthesizer using GPT-4o evaluation (Section III-C1), where step clusters were manually tested for bug reproduction and categorized. The findings indicate 66.28% of clusters were ”True” (successful reproduction), while 33.72% were ”Faulty” (preventing reproduction). The primary reasons for faulty plans were ”Wrong Command” (48.28%), ”Missing Step” (34.49%), and ”Logic Error” (31.03%). After our thorough manual analysis, 14 reports were classified as ”Irreproducible” due to factors like hardware issues or original report errors, not synthesizer faults.
IV-B1 Inter-Rater Agreement for Step Synthesizer